A commercial sample of hydrogen peroxide is labelled as 10 volume. Its percentage strength is nearly
A. 1% B. 3% C. 10% D. 90%

Grade:Upto college level

2 Answers

Avinash
8 years ago
C. 10%
Dishika Rastogi
6 years ago
The answer is 3%.
Volume Strength= (68/22.4)×V [2H2O2= O2+2H2O
Now 34×2 grams of Hydrogen Peroxide gives 22.4 litres Oxygen.
This implies that V volume of O2 reacts with (68/22.4)×V H2O2]
Put V=10 and convert it into %. You will get 3%.
I hope this helps.